OpenStudy (anonymous):

Mary invested one amount at 9% simple interest, and a second amount at 6% interest, earning $18.12 in one year. If she had switched the amounts, she would have earned $17.88. What were the two amounts?

OpenStudy (anonymous):

Let x is the first amount, y is the second: then 9x + 6y = 18.12 Can you do the second one?

OpenStudy (anonymous):

would it be x + y = 17.88

OpenStudy (anonymous):

SWITCH the amounts x, y: 9x + 6y

OpenStudy (anonymous):

6x + 9y = 17.88

OpenStudy (anonymous):

do you add the two together or just solve for the two equations

OpenStudy (anonymous):

I am a little lost would the first answer for the first equation be x = 2.013333

OpenStudy (anonymous):

not sure what you are saying. can you explain one of the equations

OpenStudy (anonymous):

9x + 6y = 18.12 (1) -> 6y = 18.12 - 9x Do you understand this part?

OpenStudy (anonymous):

yes does that apply to the other equation also

OpenStudy (anonymous):

ok give me a min to try to figure it out please

OpenStudy (anonymous):

6y = 18.12 - 9x divide all by 6 y=3.02 - 9x/6 this is what i am getting and i dont think it is right

OpenStudy (anonymous):

6x + 9y = 17.88 6x + 9(3.02 - 9x/6) = 17.88 6x + 27.18 - 81x/54 = 17.88 27.18 - 75x/54 = 17.88 -75x/54 = -9.3 -75x = -502.2 x = 6.696

OpenStudy (anonymous):

do you not undersand what i am saying or is it the wrong way

OpenStudy (anonymous):

ok ty

OpenStudy (anonymous):

6x + 9(3.02 - 9x/6) = 17.88 6x + 27.18 - 13.5x = 17.88 Question?

OpenStudy (anonymous):

oh i see how you did that

OpenStudy (anonymous):

9x/ 6 = 1.5x

OpenStudy (anonymous):

6x + 27.18 -13.5x = 17.88 7.5x + 27.18 = 17.88 7.5x = 45.06 x=6.008

OpenStudy (anonymous):

6x + 27.18 -13.5x = 17.88 -> - 7.5x = 17.88 - 27.18

OpenStudy (anonymous):

-7.5x = 17.88 - 27.18 -7.5 =-9.3

OpenStudy (anonymous):

x= 1.32

OpenStudy (anonymous):

x = 1.32

OpenStudy (anonymous):

x = 1.24 Do you use calculator?

OpenStudy (anonymous):

yes sorry must hit wrong number

OpenStudy (anonymous):

i am going to try

OpenStudy (anonymous):

9x + 6y = 18.12 this equation right

OpenStudy (anonymous):

9x + 6y = 18.12 9(1.24) + 6y = 18.12 11.16 + 6y = 18.12 6y = 6.96 y = 1.16

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

no ty for all your help and being patient with me

OpenStudy (anonymous):

i am still trying to learn the elimination method

OpenStudy (anonymous):

yes i do can i give you a problem

OpenStudy (anonymous):

9x + 6y = 18.12, 6x + 9y = 17.88

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

ause i get confused

OpenStudy (anonymous):

not sure what you mean. like this 6(9x + 6y = 18.12) 54x + 36y = 108.72

OpenStudy (anonymous):

ok then what

OpenStudy (anonymous):

6x + 9y = 17.88 9(6x + 9y = 17.88) 54x + 81y = 160.92

OpenStudy (anonymous):

no

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

45y = 52.20 y = 1.16

OpenStudy (anonymous):

So now, it's getting so easy! I've guided you through both methods. Next time it's up to you to use which one :)

OpenStudy (anonymous):

ty again

OpenStudy (anonymous):

Don't forget to click on the blue "Best Response" button if you're happy with the guidance ;)